Here I go again: Please show me how to find the slope and y-intercept of the line y = 4/3x - 3. Thanks
Found 2 solutions by Alan3354, texttutoring:
Answer by Alan3354(69443)

Please show me how to find the slope and y-intercept of the line y = 4/3x - 3
--------------
It's in slope-intercept form now. y = mx + b
m is the slope = 4/3
b = -3

Answer by texttutoring(324)

The equation of a line is y=mx+b, where m=slope and b=y-intercept

In your equation, m=4/3 and b=-3

The slope is 4/3 and the y-intercept is -3.